# Derivative-Aligned Anticipation of Forbush Decreases from Entropy and Fractal Markers

### v1

@@ 3 Methods
Window length (min) $W$ 130 $\{90,\,130,\,180\}$
Stride (min) $\Delta$ 1 $\{1,\,5\}$
@@
Persistence (min) $d$ 5 $\{5,\,10\}$
Derivative winsorization 1–99% off / 1–99%
We also list the default parameterization of each invariant estimator (See Table 2) as applied to the one–minute sliding windows. These settings follow standard practice in the literature and were validated through small–scale sensitivity sweeps to ensure stable behavior around FD onsets.
Table 2: Default parameters for invariant estimators (per sliding window).
Invariant Key parameters Default
Permutation entropy Embedding $(m,d)$ m=3m{=}3, d=1d{=}1
Sample / Approx. entropy $(m,r)$ m=2m{=}2, $r{=}0.2\,\sigma$
Spectral entropy Spectrum Normalized periodogram
Shannon entropy Binning Freedman–Diaconis (auto)
Lempel–Ziv complexity Symbolization Median threshold (binary)
Hurst (DFA) Orders / scales DFA(1–2), scales [8,W/2][8,\,W/2] min
Higuchi fractal dim. $k_{\max}$ 1010
Katz / Petrosian FD Path def. Standard definitions
Correlation dimension Embedding $(m,d)$ AMI first min.; FNN stop, $m\!\leq\!8$
Correlation dimension Scaling fit Linear region of $\log C(r)$ vs. $\log r$
@@ 4 Results and Evaluation
We applied the pipeline of §3 to the three FD episodes using minute–resolution NMDB series. Figure 1 illustrates representative count records with onset brackets. All markers were computed as sliding-window invariants and analyzed through their EWM-smoothed derivatives. Unless otherwise noted, negative lags indicate that the marker anticipates the derivative minimum of the raw count.
Across the three events, multiple invariants depart from their pre–onset baselines before the count derivative reaches its most negative excursion. Entropy–type measures (permutation, sample, spectral) typically trend downward ahead of onset, consistent with increased regularity prior to the sharp drop; algorithmic complexity (Lempel–Ziv) exhibits anticipatory dips in most stations; and geometric/fractal descriptors (Higuchi, Katz, Petrosian) show reductions compatible with a temporary smoothing of trajectories. DFA–based Hurst exponents often decrease into the onset, indicating a transient weakening of long–range dependence. The amplitude and persistence of these patterns are station–dependent and largest for the extreme 2024-05-10 event; the rapid 2024-03-24 case shows the same signatures compressed in time, in line with the fast decrease–recovery morphology and strong anisotropy reported for that storm (Mavromichalaki et al., 2024; Papailiou et al., 2024).
Table 3 reports the global ranking across stations and events under the operational detector (§3). Several invariants exhibit large negative median leads (hundreds of minutes) with substantial station coverage. In particular, Lempel–Ziv, sample entropy, Katz fractal dimension, and Shannon entropy occupy the top tier, combining large anticipatory medians with high fractions of flagged stations. Results are expressed in minutes (negative = precedes) and the “Sig. stations [%]” column reflects the percentage of stations that show sustained pre–onset excursions according to the robust threshold–and–persistence rule (no cross–correlation nor FDR is applied).
Table 3: Global ranking of invariants by robustness score $R_{k}$ and median lead $\widetilde{\ell}_{k}$ (min; negative = precedes).
Invariant $\widetilde{\ell}_{k}$ IQR Sig. stations [%] $R_{k}$
Lempel–Ziv -461.5 61.2 88 0.85
Sample entropy -401.5 80.8 78 0.80
Approx. entropy -376.0 225.5 41 0.70
Correlation dimension -398.0 160.8 75 0.70
DFA (Hurst) -361.0 258.0 59 0.70
Higuchi fractal dim. -290.5 203.5 97 0.70
Hurst -318.5 224.0 81 0.70
Katz fractal dim. -441.0 169.2 81 0.70
MF–Hurst (band b) -375.0 130.8 94 0.70
Permutation entropy -312.0 218.5 72 0.70
Petrosian fractal dim. -215.0 253.0 53 0.70
Shannon entropy -399.5 124.0 100 0.70
Spectral entropy -91.0 164.2 97 0.60
Entropy -63.0 95.2 100 0.57
Figure 2 visualizes the median lead per invariant and event. Measures tied to long-range dependence and roughness (Hurst) and to graph-based geometry (Katz) are consistently anticipatory, while permutation entropy is event-dependent: it lags for April 2023 and March 2024, but turns clearly anticipatory in the extreme May 2024 case.
Figure 2: Median lead by invariant and event (min; negative = precedes).
Table 4 summarizes per-event medians and station-level significance. The three episodes display distinct morphologies.
Table 4: Per–event summary: median lead $\widetilde{\ell}_{k}$ (min; negative = precedes) and percent of stations with significant pre–onset change.
2023-04-23 2024-03-24 2024-05-10
Invariant $\widetilde{\ell}_{k}$ Sig. [%] $\widetilde{\ell}_{k}$ Sig. [%] $\widetilde{\ell}_{k}$ Sig. [%]
app_entropy -369.0 0 -278.0 100 -395.0 18
corr_dim -394.5 50 -311.0 100 -418.0 73
dfa -414.0 40 -259.0 100 -376.0 36
entropy -96.5 100 -69.0 100 -25.0 100
higuchi_fd -408.0 90 -277.0 100 -286.0 100
hurst -376.5 60 -257.0 100 -328.0 82
katz_fd -463.0 90 -257.0 100 -440.0 55
lepel_ziv -445.5 90 -450.0 100 -476.0 73
mfhurst_b -404.0 90 -319.0 100 -404.0 91
permutation_entropy -321.0 70 -293.0 100 -319.0 45
petrosian_fd -186.0 70 -220.0 36 -238.0 55
sampen -421.0 60 -357.0 100 -421.0 73
shannon_entropy -439.0 100 -250.0 100 -423.0 100
spectral_entropy -84.0 100 -89.0 91 -116.0 100
For 2024-05-10 (extreme FD), most invariants exhibit broad negative leads with high station coverage: Shannon entropy (423-423 min; $100\%$), multifractal Hurst band $b$ (404-404 min; $91\%$), Lempel–Ziv (476-476 min; $73\%$), sample entropy (421-421 min; $73\%$), and Katz fractal dimension (440-440 min; $55\%$). Correlation dimension and Higuchi also anticipate with substantial coverage (418-418 and 286-286 min; $73\%$ and $100\%$, respectively). This event concentrates the strongest, most coherent anticipatory signals across the network. The corresponding station-wise lag distributions are shown in Figure 3 (upper-left panel).
For 2024-03-24 (rapid FD), precedence is widespread but more compressed. Many measures remain clearly anticipatory with near-universal coverage: Shannon entropy (250-250 min; $100\%$), sample entropy (357-357 min; $100\%$), Lempel–Ziv (450-450 min; $100\%$), multifractal Hurst $b$ (319-319 min; $100\%$), and Higuchi (277-277 min; $100\%$). Spectral entropy also anticipates (89-89 min; $91\%$). These patterns match the short decrease–recovery morphology reported for this storm, with leads present but temporally compact. Station-wise distributions are shown in Figure 3 (upper-right panel).
For 2023-04-23, anticipatory behavior is mixed but still present across several families. Katz (463-463 min; $90\%$), Shannon (439-439 min; $100\%$), Lempel–Ziv (445.5-445.5 min; $90\%$), sample entropy (421-421 min; $60\%$), multifractal Hurst $b$ (404-404 min; $90\%$), and Higuchi (408-408 min; $90\%$) show robust negative medians with sizable station coverage, while Petrosian exhibits more modest anticipation (186-186 min; $70\%$). Station-wise distributions are shown in Figure 3 (bottom panel).
Figure 3: Station-wise lag distributions by invariant for three events studied (negative = precedes).
A compact operational panel emerges across events: Katz fractal dimension, DFA–based Hurst, Shannon and sample entropy, and Lempel–Ziv combine large negative medians (often $\mathcal{O}(10^{2}\text{\textendash}10^{3})$ min) with high station coverage under the operational detector. This panel yields substantial lead for the extreme case (2024-05-10) and retains predictive value for the rapid morphology of 2024-03-24.
The three overlays in Fig. 4 illustrate the typical temporal organization of the marker derivatives relative to the count derivative for the selected stations. For the extreme FD on 2024-05-10 (CALG), the count derivative shows a broad, persistent negative episode. All four markers display sustained departures before and during the drop, with Hurst exhibiting the earliest and largest negative swings and Katz and Shannon remaining below baseline for extended periods. Sample entropy rises and then turns negative as the decrease develops, consistent with the strong anticipatory medians reported for this event.
For the rapid FD on 2024-03-24 (LMKS), the onset is sharp and the recovery compressed. Marker excursions are shorter and partially contemporaneous: Hurst shows bursty oscillations with brief negative lobes just before onset, while Katz and Shannon remain closer to zero with limited lead. This visual compression matches the mixed, sometimes positive, event-wise medians of entropy-like descriptors and the reduced leads of scaling-based markers.
For the 2023-04-23 event (NEWK), behavior is mixed. Katz tracks a slow downward drift that begins before the count derivative reaches its most negative values, whereas Hurst is more volatile, with pre-onset spikes and a broader envelope around the transition. Shannon entropy maintains a gentle pre-onset ridge and sample entropy exhibits wide fluctuations that straddle the onset. These patterns are consistent with the per-event summary, where several ordinal/fractal measures tend to lag, while Katz, Hurst, and Shannon retain moderate negative medians.
Overall, the overlays corroborate the heatmap and violin summaries: the May 2024 episode concentrates the strongest, most coherent anticipatory signatures across markers; the March 2024 episode compresses leads and shifts some information-theoretic measures to contemporaneous or lagging responses; and the April 2023 episode exhibits a mixed morphology with geometric and scaling markers more consistently leading than ordinal measures. The timing of these excursions is commensurate with the analysis window ($W=130$ min) and the pre–onset search bound ($L=480$ min), supporting derivative–based alignment as the operational basis for early–warning markers.
Figure 4: Derivative overlays for a fixed panel of markers (Katz fractal dimension, Hurst via DFA, Shannon entropy, sample entropy) against the EWM–smoothed derivative of the original count (native units). Negative excursions of marker derivatives ahead of the most negative count–derivative episode indicate anticipatory behavior.
@@ 5 Discussion
The analyses reveal a consistent picture in which markers tied to roughness, memory, and graph geometry tend to anticipate FD onsets, whereas some ordinal and frequency–domain measures are more event-dependent (Gil et al., 2021; Koikkalainen et al., 2025; Raath et al., 2022). The strongest and most coherent precedence arises in the extreme 2024–05–10 episode; the May 10, 2024, event reached exceptional magnitude ($\approx$15.7% at 10 GV) and was associated with GLE 74 under severe magnetospheric conditions (Abunina et al., 2025; Papaioannou et al., 2025). Hurst shows large negative excursions well before the count-derivative minimum, while Katz fractal dimension and Shannon entropy remain persistently depressed across the onset. This suggests a preconditioning phase where transport becomes less persistent and the time–series graph smooths prior to the main decrease. In contrast, the rapid 2024–03–24 case compresses or erodes these leads, indicating that when the heliospheric disturbance evolves quickly the system affords less time for precursory reorganization to develop or be captured within fixed windows; spatial coherence and anisotropy patterns reported for severe FDs are consistent with this picture (Riggi et al., 2025). The April–2023 event sits between these extremes and highlights cross–family diversity: geometric and scaling markers provide moderate lead while several ordinal descriptors lag, emphasizing that “one size fits all” is unlikely for FD morphology (Koikkalainen et al., 2025; Raath et al., 2022).
Operationally, two aspects matter. First, many markers deliver actionable lead times on the order of tens to a few hundred minutes, with high geographical coherence (Riggi et al., 2025). This favors station-agnostic deployment and ensemble decision rules. Second, the derivative-based alignment together with an operational detector (robust $z$ on pre-onset windows with bilateral threshold and persistence) sharpens transitions and reduces sensitivity to slow drifts, yielding clearer timing summaries (median leads in minutes and station coverage) without relying on cross-correlation or hypothesis testing. The heatmaps, violins, and overlays jointly argue for a compact nowcasting panel comprising Hurst (DFA), Katz FD, Shannon/Sample entropy, and Lempel–Ziv—measures that are interpretable, computationally light, and available in robust open implementations (Gil et al., 2021).
There are limitations. We use NMDB one–minute counts in station–specific native units as provided by the data source; we do not perform cutoff–rigidity homogenization or additional detector–response corrections beyond what is standard in the provider. Windowing choices ($W,\Delta,\alpha$) necessarily trade temporal precision for noise suppression; while our sensitivity sweeps indicate qualitative stability, truly optimal parameters may vary with event class and cadence. Station–level onset is defined as the time of the minimum of the first derivative of the smoothed count series; absolute timings can shift by minutes under alternative onset models. The operational detector depends on ($L,Z_{0},d$); we explored representative settings, but adaptive thresholds or per–invariant directionality may further refine detection under nonstationary turbulence. For scaling/memory estimation, DFA remains effective at minute cadence, yet recent advances show Bayesian or likelihood–based Hurst inference can improve small–sample performance at higher computational cost (Borin, 2024; Mangalam et al., 2023).
Several extensions follow naturally. Adaptive or multiscale windows could capture both gradual precursors and abrupt onsets; multiscale variants of Katz FD are a natural avenue to enhance sensitivity without prohibitive cost (Li et al., 2023). Combining invariant markers with upstream solar–wind and IMF measurements may increase lead and confidence via multimodal fusion; recent entropy–complexity analyses in solar–wind streams and ICME sheaths support the diagnostic value of information–theoretic features for regime discrimination (Koikkalainen et al., 2025). Online probabilistic scoring (e.g., Bayesian filtering or conformal risk control) would turn marker excursions into calibrated alerts. Finally, systematic evaluation across solar–cycle phases and cutoff–rigidity bands would test generalization beyond the three case studies and clarify anisotropy–related effects.

### v2

@@ 3.6 Sensitivity analysis
Window length (min) $W$ 120 $\{90,\,120,\,180\}$
Stride (min) $\Delta t$ 5 $\{1,\,5\}$
@@
Persistence (steps) $d$ 5 $\{3,\,5,\,10\}$
Min valid fraction per window 0.95 $\{0.90,\,0.95\}$
@@ 4 Results and Evaluation
We apply the pipeline described in §3 to two Forbush decrease (FD) episodes with uniform 1-minute cadence in the NMDB station network: 2023-04-23 and 2024-05-10. For each event we analyze a 28-station subset and compute sliding-window invariants using a causal window ($W=120$ min) evaluated every $\Delta t=5$ min, so that each marker value $M_{k}(t)$ depends only on measurements up to time $t$ (no look-ahead). Throughout this section, the alignment time $t_{0}$ is defined operationally as the timestamp where the exponentially weighted moving-average (EWM) first difference of the smoothed count rate reaches its minimum at each station. Marker “leads” are reported in minutes as $\ell^{*}=t_{k}-t_{0}$, where $t_{k}$ is the time of the most negative excursion of the marker derivative $\dot{M}_{k}(t)$ within the pre-onset search window; hence $\ell^{*}<0$ indicates that the marker change precedes $t_{0}$.
Onset alignment diagnostic.
Figure 2 illustrates the onset diagnostic on representative stations for each event, overlaying the raw count, the EWM-smoothed count, and the EWM-smoothed first difference. The selected $t_{0}$ consistently falls within the transition into the main decrease phase, at or near the strongest negative excursion of the smoothed first difference. We emphasize that $t_{0}$ is used here as an alignment reference anchored in the count-rate dynamics, rather than as a threshold-based physical onset; this choice enables a station-by-station synchronization of heterogeneous profiles while preserving a fully automated definition that can be applied uniformly across stations and events.
Figure 2: Onset diagnostics for representative stations and events. For each panel, the raw NM count rate (light blue) is smoothed with an exponentially weighted moving average (EWM; orange), and the EWM-smoothed first difference $\Delta x_{s}(t)$ (dark blue, right axis) is used to define the reference onset time $t_{0}$ (vertical dashed line) as described in §3. This diagnostic illustrates the event morphology and the operational alignment used to measure marker leads/lags.
A compact CORE panel.
To reduce redundancy across closely related invariants, we focus the main evaluation on a compact CORE set spanning complementary families: information-theoretic measures (Shannon, approximate, sample, and spectral entropy), geometric/roughness descriptors (Higuchi and Katz fractal dimension), and complexity/embedding indicators (Lempel–Ziv and correlation dimension). Table 2 reports pooled station-wise performance across both events, summarizing the fraction of station–event pairs flagged by the operational detector (Detect %), and the distribution of station-wise leads. Across the two events, detection rates are consistently high (78–93%), indicating that the CORE invariants frequently exhibit sustained pre-$t_{0}$ excursions under the robust threshold-and-persistence rule. Median leads are systematically negative and typically of order several hours: for example, spectral entropy shows a pooled median lead of 365.5-365.5 min (Detect 83.6%), while Lempel–Ziv and Shannon entropy yield pooled medians of 320-320 min (Detect 81.8%) and 300-300 min (Detect 92.7%), respectively. These summaries motivate the CORE panel as a parsimonious operational set that retains the strongest and most repeatable anticipatory signatures while markedly reducing the methodological density of the full metric suite.
Table 2: CORE-panel station-wise lead statistics pooled across the two events (negative lead = precedes onset). Detect (%) is the fraction of station–event pairs flagged by the operational detector. NN is the number of station–event pairs with a valid lead estimate for that invariant (out of 56 total pairs; 28 stations $\times$ 2 events).
Invariant Detect (%) N Median lead (min) IQR (min) Q25 Q75
shannon_entropy 92.7 51 -300.0 204.5 -401.0 -196.5
app_entropy 90.9 50 -203.0 197.0 -308.5 -111.5
higuchi_fd 87.3 48 -296.0 175.5 -367.8 -192.2
sampen 85.5 47 -273.0 218.5 -388.5 -170.0
corr_dim 85.5 47 -227.0 221.0 -340.0 -119.0
spectral_entropy 83.6 46 -365.5 153.2 -424.8 -271.5
lepel_ziv 81.8 45 -320.0 187.0 -432.0 -245.0
katz_fd 78.2 43 -305.0 282.0 -407.5 -125.5
Detection coverage across invariants.
Figure 3 summarizes the detection coverage of each CORE invariant under the same robust threshold-and-persistence rule. The ordering is consistent with Table 2: most invariants trigger in a large majority of station–event pairs, indicating that the anticipatory signal is not confined to a small subset of stations. Importantly, this figure isolates coverage from timing: a metric may be highly detectable yet exhibit a shorter median lead, so Detect (%) should be interpreted jointly with the lead distributions (Figs. 45).
Figure 3: Detection rate by invariant for the CORE panel under the operational robust threshold-and-persistence detector (§3). Bars report the fraction of stations that exhibit sustained pre-onset excursions according to the predefined significance rule.
Station-wise lead dispersion and event dependence.
Figure 4 shows the station-wise lead distributions for the CORE panel separately for each event using median ticks and interquartile ranges (IQR). The distributions reveal substantial station-to-station variability, with IQRs commonly spanning $\sim$2–4 hours depending on the invariant and event. This dispersion is expected given differences in geomagnetic cutoff rigidity, asymptotic directions, and local count-rate noise characteristics, and it reinforces the need for robust, station-level criteria rather than single-station inference.
Figure 4: Station-wise lag distributions for the CORE marker panel, shown separately for the two analyzed events. For each invariant, the horizontal segment summarizes the cross-station spread of the estimated lag $\ell^{\*}$ (in minutes) relative to the reference time $t_{0}$; negative values indicate anticipation (marker changes precede $t_{0}$). The vertical tick marks the station-level median lag for each invariant, and the dashed vertical line at 0 marks contemporaneous response.
Complementarily, Figure 5 overlays boxplots and individual station points for both events, making explicit both the central tendency and the prevalence of outliers (e.g., stations with near-zero or weakly negative leads) that contribute to the observed IQR.
Figure 5: Lead (lag) distributions by CORE invariant and event across the station network. Boxplots summarize the station-wise lags $\ell^{\*}$ (min) for each metric; dots show individual-station values. Negative values indicate anticipation (marker precedes $t_{0}$), and the horizontal dashed line at 0 marks contemporaneous response. This view highlights both robustness (tight interquartile ranges) and heterogeneity (tails/outliers) across stations.
Event-wise summaries.
Table 3 reports per-event CORE summaries (Detect %, median lead, and IQR). For 2023-04-23, Shannon entropy exhibits a strong median lead of 358-358 min with 92.6% detection, accompanied by spectral entropy at 347-347 min (77.8%) and both Higuchi fractal dimension and Lempel–Ziv near 308-308 min (81.5%). The 2024-05-10 episode displays a similar anticipatory pattern but with shifts in which invariants lead most strongly: spectral entropy reaches a median of 383-383 min (89.3%), Lempel–Ziv 343-343 min (82.1%), and sample entropy 339-339 min (89.3%), while Shannon entropy remains highly detectable (92.9%) with a more moderate median lead (259-259 min). These differences are consistent with the fact that the events exhibit distinct decrease morphologies and station responses, yet the CORE panel maintains broad pre-$t_{0}$ detectability and predominantly negative lead distributions in both cases.
Table 3: CORE-panel performance by event (station-wise). Detect (%) is the fraction of stations flagged by the operational detector; leads are in minutes (negative = precedes $t_{0}$).
2023-04-23 2024-05-10
Invariant Detect (%) Median IQR Detect (%) Median IQR
shannon_entropy 92.6 -358.0 216.0 92.9 -259.0 175.0
app_entropy 88.9 -207.0 148.8 92.9 -201.0 245.8
higuchi_fd 81.5 -308.5 178.5 92.9 -279.5 174.2
sampen 81.5 -219.0 120.0 89.3 -339.0 201.0
corr_dim 85.2 -215.0 238.0 85.7 -231.5 200.0
spectral_entropy 77.8 -347.0 135.0 89.3 -383.0 158.0
lepel_ziv 81.5 -308.0 142.8 82.1 -343.0 198.0
katz_fd 77.8 -261.0 259.0 78.6 -321.0 238.5
Qualitative behavior of markers and their derivatives.
To connect the summary statistics with the underlying time-domain behavior, Figure 6 and 7 present representative time series of $M_{k}(t)$ and $\dot{M}_{k}(t)$ for selected stations in each event. In the 2024-05-10 case, multiple CORE markers display coherent pre-$t_{0}$ structure: entropy-type measures and complexity/embedding indicators tend to decline into the decrease phase, while geometric descriptors show concomitant changes consistent with a temporary reduction in irregularity. The marker derivatives $\dot{M}_{k}(t)$ exhibit sustained negative excursions in the pre-$t_{0}$ window, providing the operational basis for early flagging. For 2023-04-23 the same qualitative organization is present but is more heterogeneous across stations, which is reflected in broader IQRs for several invariants in Table 3 and Figure 5. Importantly, these examples supply the visual grounding requested by the referee: they demonstrate that the proposed lead extraction is tied to identifiable features in the marker trajectories rather than being an artifact of tabulated summaries.
Figure 6: CORE marker time series $M_{k}(t)$ (left column) and their EWM-smoothed derivatives $\dot{M}_{k}(t)$ (right column) for two representative stations during the 2024-05-10 event. The vertical dashed line marks $t_{0}$. The computation uses a causal sliding window (step and window length as annotated in the panels).
Figure 7: Same as Fig. 6, but for the 2023-04-23 event. These examples illustrate how anticipatory excursions appear as sustained departures in $\dot{M}_{k}(t)$ prior to $t_{0}$, while also showing station-dependent variability.
Summary.
Overall, the evaluation indicates that a compact CORE panel can provide reproducible, station-wise anticipatory signatures relative to the operational alignment time $t_{0}$, with median leads typically ranging from $\sim$3 to 6 hours and detection rates above $\sim$80% for most CORE invariants. The combination of onset diagnostics (Figure 2), distributional summaries (Table 2, Figure 4, Figure 5), and representative marker trajectories (Figure 6 and 7) yields a complete and interpretable account of both the quantitative performance and the qualitative mechanisms underlying the anticipatory behavior.
@@ 5 Discussion
Across two FD episodes and a 28-station network subset, the results support a consistent operational picture: a compact set of information-theoretic, geometric, and complexity/embedding invariants exhibits sustained pre-$t_{0}$ structure with high station-wise detection coverage and predominantly negative lags. In particular, entropy-type measures (Shannon, spectral, sample, approximate), geometric descriptors (Katz and Higuchi fractal dimensions), and complexity/embedding indicators (Lempel–Ziv and correlation dimension) form a coherent panel that repeatedly departs from pre-onset baselines before the steepest decline in the smoothed count-rate dynamics. The fact that multiple families respond in a temporally ordered way, rather than a single metric dominating, mirrors broader findings that entropy/complexity and geometric signatures can capture regime shifts in space-physics time series and related geospace observables (Koikkalainen et al., 2025; Raath et al., 2022; Gil et al., 2021).
A key implication is that the anticipatory behavior appears as a network signature rather than a single-station artifact: detection fractions remain high across stations, while lead times exhibit substantial dispersion. Such heterogeneity is expected given differences in cutoff rigidity, asymptotic directions, and local noise characteristics, and is consistent with prior evidence that FD development and anisotropy can vary substantially across latitude bands and detector types during severe disturbances (Riggi et al., 2025). Accordingly, the most robust operational use case is station-agnostic deployment with ensemble aggregation (e.g., requiring agreement across multiple stations and/or multiple markers) rather than relying on any individual station as definitive.
Event dependence is also informative. The 2024–05–10 episode shows particularly coherent anticipation across the CORE panel, consistent with its classification as an extreme heliospheric disturbance. Reports on the May 2024 storm period emphasize unusually strong FD behavior and associated high-energy particle activity (including the GLE 74 episode), providing context for why multiple statistical families might exhibit sustained departures over extended pre-$t_{0}$ intervals (Abunina et al., 2025; Papaioannou et al., 2025). In contrast, the 2023–04–23 event displays the same qualitative organization but with broader station-to-station spreads, reinforcing that FD morphology and observational geometry modulate both the timing and the clarity of precursory signatures (Riggi et al., 2025). Importantly, our interpretation remains operational: we do not claim that any particular invariant uniquely identifies a physical microprocess; rather, we show that several complementary invariants reliably change before the steepest count-rate decline, which is the key requirement for an early-flagging scheme.
Methodologically, two design choices matter for how these results should be read. First, $t_{0}$ is used as an alignment reference derived from the count-rate dynamics (minimum of the EWM-smoothed first difference), not as a universal, threshold-based physical onset. This enables station-by-station synchronization of heterogeneous profiles without event-specific tuning. Second, the marker computation uses a causal window, which avoids look-ahead and makes the extracted leads compatible with online deployment. In this setting, the robust threshold-and-persistence detector provides a practical way to translate marker derivatives into reproducible timing summaries (coverage, medians, IQRs) without cross-correlation or formal hypothesis testing, while preserving interpretability of the underlying marker trajectories (Koikkalainen et al., 2025; Raath et al., 2022).
Several limitations remain. We use NMDB one-minute counts in native station units and do not attempt rigidity homogenization or additional response corrections beyond the data provider. Window length, step, and smoothing parameters necessarily trade temporal precision for noise suppression, and truly optimal settings may vary across event classes and station noise regimes. Moreover, although geometric and entropy/complexity markers are computationally light, some estimators (especially those tied to memory/scaling in broader screenings) can be sensitive in finite samples; recent work suggests that Bayesian or likelihood-based approaches to long-memory characterization may improve robustness at additional computational cost (Borin, 2024; Mangalam et al., 2023). Finally, while our results are consistent across two contrasting events, generalization requires larger multi-event evaluation across solar-cycle phases and rigidity bands, particularly under strong anisotropy.
Natural extensions follow. Multi-scale or adaptive windows could better accommodate both gradual precursors and abrupt transitions, and directionality could be handled explicitly for invariants whose excursions are not consistently signed across all morphologies. Multiscale generalizations of geometric descriptors (e.g., Katz-type constructions) are a promising route to increase sensitivity without prohibitive cost (Li et al., 2023). Beyond univariate NM counts, combining the invariant panel with upstream solar-wind/IMF context could provide multimodal confirmation and improve alert confidence; recent entropy/complexity analyses of solar-wind structures support the diagnostic value of information-theoretic features for regime discrimination (Koikkalainen et al., 2025). These steps would move the present case-study evidence toward calibrated, operational early-flagging in a network setting.